# Unlearning Concepts in Diffusion Model via Concept Domain Correction and Concept Preserving Gradient

### v2

@@ Introduction
A naive approach is to filter out sensitive information and then retrain the entire model from scratch (Rombach et al. 2022; Rao 2023; Kim et al. 2025). However, given that these generative models are typically trained on large-scale datasets (Schuhmann et al. 2022), this method can be prohibitively costly. Another strategy involves using a Safety Checker to detect the presence of inappropriate content in generated images (Rombach et al. 2022). However, this approach relies heavily on the performance of the detector and is limited by the biases.
In addressing these challenges, Machine Unlearning (MU) (Xu et al. 2023; Gao et al. 2024b, a) has emerged as a potentially promising solution. Fundamentally, MU is a method that enables models to unlearn the memory of sensitive information contained within their training datasets, thereby preventing the generation of images involving sensitive concepts through crafted prompts. Recently, a series of studies have been introduced that focus on the unlearning concepts in diffusion models (Gandikota et al. 2023; Zhang et al. 2023; Gandikota et al. 2024; Kumari et al. 2023; Lyu et al. 2023; Zhang et al. 2024c; Guo, Wu, and Wan 2024; Liu et al. 2023). For instance, ESD (Gandikota et al. 2023) fine-tunes the model to predict in the opposite direction of classifier-free guidance, effectively aligning the targeted concept with that of an empty string. ConAbl (Kumari et al. 2023) proposes linking a target concept to a predefined anchor concept by minimizing the L2 distance between the predicted noises for these two concepts. SPM (Lyu et al. 2023) employs a Latent Anchor approach to eliminate concepts and uses a similarity-based retention loss to preferentially weight surrogate concepts, thereby preserving non-target concepts.
@@
Text-to-image diffusion models (Rombach et al. 2022; Saharia et al. 2022; Nichol et al. 2022; Zhang et al. 2024b; Cheng et al. 2024; Yang et al. 2025; Singh et al. 2024) have recently shown impressive potential in generating high-quality images, typically trained on large-scale, web-crawled datasets such as LAION-5B (Schuhmann et al. 2022). However, these uncurated datasets often include harmful or copyrighted content, which can lead the models to generate sensitive or Not Safe For Work (NSFW) images. The research community has initiated various efforts to address this issue, which can generally be classified into three main categories: dataset filtering, model fine-tuning, and post-generation classification.
@@
Model fine-tuning methods focus on adjusting the model’s weights to prevent the generation of harmful content (Zhang et al. 2025, 2024a; Pan et al. ; Park et al. 2024; Ko et al. 2024; Kim, Min, and Yang 2025). ESD (Gandikota et al. 2023) aligns the probability distributions of the targeted concept with that of an empty string. However, this technique may lead to a collapse problem due to the lack of explicit control. Forget-Me-Not (Zhang et al. 2023) employs attention re-steering to identify and modify attention maps associated with specific concepts within the cross-attention layers of the diffusion U-Net architecture. However, both methods lack a mechanism for preserving the integrity of other non-target concepts. UCE (Gandikota et al. 2024) uses a closed-form solution to adjust cross-attention weights, recalibrating them to induce deliberate changes in the keys and values associated with specific text embeddings for the concepts to be edited, while aiming to limit the impact on a separate set of concepts to be retained. ConAbl (Kumari et al. 2023) suggests associating a target concept with a predefined anchor concept by minimizing the L2 distance between the predicted noises for these two concepts, and introduces a regularization loss to preserve the integrity of the anchor concept. However, these approaches do not fully resolve the tension between unlearning and retention objectives, potentially impacting the optimization efficiency of both goals. SPM (Lyu et al. 2023) introduces a novel Latent Anchoring fine-tuning strategy combined with a similarity-based retention loss to differentially weight surrogate concepts that are distant in the latent space. While this approach can alleviate the contradictions between the two objectives, it does not directly resolve their conflicts. Our method directly employs a gradient surgery technique to manipulate the gradients of the two conflicting objective functions. Furthermore, previous approaches heavily depend on the configuration of the training set, such as forming pairs of anchor and target concepts or directly manipulating concepts, which can lead to decreased generalization performance when applied beyond the training scope. In contrast, our method modifies the concept at the distribution level, effectively mitigating this problem.
@@
Figure 4: The I2P benchmark results using NudeNet Detection. We use the generations produced by Stable Diffusion v1.4 as baseline. The bin plots illustrate the reduction in explicit content through various unlearning methods for the nudity concept.
@@
Figure 5: Visualization examples of artistic styles unlearning. Upper: “Van Gogh”. lower: “Picasso”.
@@
We utilize the CLIP Score (CS), CLIP Accuracy (CA), and the FID as metrics to evaluate the performance on instance/style unlearning (Hessel et al. 2021; Kumari et al. 2023; Heusel et al. 2017; Li et al. 2024a; Liu et al. 2024; Li et al. 2024b). For unlearning celebrities, We use Identity Score Matching (ISM) which compute the cosine distance between the embedding of the generated face and the average face embedding of the entire clean image set (Van Le et al. 2023).
@@
Table 2: The results of unlearning celebrities.
@@
Table 3: Ablation study on assessing the impact of L2 loss (L2) versus concept preserving gradient (CP).
Removal of Explicit Content.
We adopt the I2P benchmark (Schramowski et al. 2023) for evaluating the performance of unlearning inappropriate content. A nudity detection model  [^1] https://github.com/notAI-tech/NudeNet/tree/v3 is applied to quantify the number of nude body parts. As depicted in Figure 4, our method manages to achieve a total reduction of 75.67%, surpassing the results obtained using ESD and ConAbl. Although it falls short of the total number reduction achieved by SPM, it holds an advantage in mitigating exposure of sensitive body parts such as “female breast” and “buttocks”.

### v3

@@ Introduction
A naive approach is to filter out sensitive information and then retrain the entire model from scratch (Rombach et al. 2022; Rao 2023; Kim et al. 2025). However, given that these generative models are typically trained on large-scale datasets (Schuhmann et al. 2022), this method can be prohibitively costly. Another strategy involves using a Safety Checker to detect the presence of inappropriate content in generated images (Rombach et al. 2022; Liu et al. 2024b; Wang et al. 2024; Gao et al. 2024c; Yoon et al. 2024). However, this approach relies heavily on the performance of the detector and is limited by the biases.
In addressing these challenges, Machine Unlearning (MU) (Xu et al. 2023; Gao et al. 2024b, a; Zhou et al. 2024) has emerged as a potentially promising solution. Fundamentally, MU is a method that enables models to unlearn the memory of sensitive information contained within their training datasets, thereby preventing the generation of images involving sensitive concepts through crafted prompts. Recently, a series of studies have been introduced that focus on the unlearning concepts in diffusion models (Gandikota et al. 2023; Zhang et al. 2023; Gandikota et al. 2024; Kumari et al. 2023; Lyu et al. 2023; Zhang et al. 2024d; Guo, Wu, and Wan 2024; Liu et al. 2023; Zhang et al. 2024a; Truong, Dang, and Le 2024). For instance, ESD (Gandikota et al. 2023) fine-tunes the model to predict in the opposite direction of classifier-free guidance, effectively aligning the targeted concept with that of an empty string. ConAbl (Kumari et al. 2023) proposes linking a target concept to a predefined anchor concept by minimizing the L2 distance between the predicted noises for these two concepts. SPM (Lyu et al. 2023) employs a Latent Anchor approach to eliminate concepts and uses a similarity-based retention loss to preferentially weight surrogate concepts, thereby preserving non-target concepts.
@@
Text-to-image diffusion models (Rombach et al. 2022; Saharia et al. 2022; Nichol et al. 2022; Zhang et al. 2024c; Cheng et al. 2024; Yang et al. 2025; Singh et al. 2024) have recently shown impressive potential in generating high-quality images, typically trained on large-scale, web-crawled datasets such as LAION-5B (Schuhmann et al. 2022). However, these uncurated datasets often include harmful or copyrighted content, which can lead the models to generate sensitive or Not Safe For Work (NSFW) images. The research community has initiated various efforts to address this issue, which can generally be classified into three main categories: dataset filtering, model fine-tuning, and post-generation classification.
@@
Model fine-tuning methods focus on adjusting the model’s weights to prevent the generation of harmful content (Zhang et al. 2025, 2024b; Park et al. 2024; Ko et al. 2024; Kim, Min, and Yang 2025). ESD (Gandikota et al. 2023) aligns the probability distributions of the targeted concept with that of an empty string. However, this technique may lead to a collapse problem due to the lack of explicit control. Forget-Me-Not (Zhang et al. 2023) employs attention re-steering to identify and modify attention maps associated with specific concepts within the cross-attention layers of the diffusion U-Net architecture. However, both methods lack a mechanism for preserving the integrity of other non-target concepts. UCE (Gandikota et al. 2024) uses a closed-form solution to adjust cross-attention weights, recalibrating them to induce deliberate changes in the keys and values associated with specific text embeddings for the concepts to be edited, while aiming to limit the impact on a separate set of concepts to be retained. ConAbl (Kumari et al. 2023) suggests associating a target concept with a predefined anchor concept by minimizing the L2 distance between the predicted noises for these two concepts, and introduces a regularization loss to preserve the integrity of the anchor concept. However, these approaches do not fully resolve the tension between unlearning and retention objectives, potentially impacting the optimization efficiency of both goals. SPM (Lyu et al. 2023) introduces a novel Latent Anchoring fine-tuning strategy combined with a similarity-based retention loss to differentially weight surrogate concepts that are distant in the latent space. While this approach can alleviate the contradictions between the two objectives, it does not directly resolve their conflicts. Our method directly employs a gradient surgery technique to manipulate the gradients of the two conflicting objective functions. Furthermore, previous approaches heavily depend on the configuration of the training set, such as forming pairs of anchor and target concepts or directly manipulating concepts, which can lead to decreased generalization performance when applied beyond the training scope. In contrast, our method modifies the concept at the distribution level, effectively mitigating this problem.
@@
Figure 4: Visualization examples of artistic styles unlearning. Left: Unlearning ”Van Gogh”. Right: Unlearning ”Picasso”. The first row represents the forgotten style, while the subsequent rows represent other non-target concepts.
@@
We utilize the CLIP Score (CS), CLIP Accuracy (CA) and FID as metrics to evaluate the performance on the unlearning of instances and styles (Hessel et al. 2021; Kumari et al. 2023; Heusel et al. 2017; Li et al. 2024a; Liu et al. 2024a; Li et al. 2024b). A concept to be evaluated is incorporated into 80 templates, with each template yielding 10 images. For unlearning celebrities, We use Identity Score Matching (ISM) which compute the cosine distance between the embedding of the generated face and the average face embedding of the entire clean image set (Van Le et al. 2023).
@@
Table 2: Unlearning Celebrities: We selected three common celebrities and unlearned two of them. Our method successfully retains the other celebrities while accurately forgetting the target concept.
@@
Table 3: Ablation Study on Assessing the Impact of L2 Loss (L2) vs. Concept Preserving Gradient (CP). It can be observed that without any retain loss, the unlearning process fails to preserve other concepts effectively. On the other hand, directly using L2 loss leads to a degradation in unlearning performance. Our proposed CP method strikes a balance between the two, achieving better results in both unlearning and concept retention.